# Testing the hypothesis of vector X17 boson by $D$ meson, Charmonium, and $\phi$ meson decays

### v1

@@ Abstract
The recent ATOMKI experiments provided evidence pointing towards the existence of an X17 boson in the anomalous nuclear transitions of Beryllium-8, Helium-4, and Carbon-12. In this work, we consider X17 boson contributions to the previously measured $D$ meson decays which include $D_{s}^{*+}\rightarrow D_{s}^{+}e^{+}e^{-}$, $D_{s}^{*+}\rightarrow D_{s}^{+}\gamma$, $D^{*0}\rightarrow D^{0}e^{+}e^{-}$, and $D^{*0}\rightarrow D^{0}\gamma$, as well as the measured decays of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, and $\phi\rightarrow\eta\gamma$. Using the data of the above meson decays, we perform a fitting to the coupling parameters $\varepsilon_{u},\varepsilon_{c}$, and $\varepsilon_{s}$ by treating the couplings $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent from each other rather than assuming the generation universality $\varepsilon_{u}=\varepsilon_{c}$. It is found that the above fitting renders $|\varepsilon_{c}|=3\times 10^{-4}$, $|\varepsilon_{s}|=4\times 10^{-5}$ and a huge magnitude for $\varepsilon_{u}$, which is in serious tension with $\left|\varepsilon_{u}\right|$ determined from ATOMKI measurements.
@@ I Introduction
The results reported in Krasznahorkay:2015iga gives $m_{X}=16.70\pm 0.35({\rm stat})\pm 0.5({\rm syst})~{}\rm{MeV/c^{2}}$ and the ratio $\Gamma(^{8}{\rm Be}(18.15)\rightarrow~{}^{8}{\rm Be}+{X})\times{\rm BR}({X}\rightarrow e^{+}e^{-})/\Gamma(^{8}{\rm Be}(18.15)\rightarrow~{}^{8}{\rm Be}+\gamma)\equiv R_{X}=5.8\times 10^{-6}$, with a statistical significance of $6.8\sigma$. Later analysis from more 8Be experiments gave the updated results, $m_{X}=17.01\pm 0.16~{}\rm{MeV/c^{2}}$ and $R_{X}=(6\pm 1)\times 10^{-6}$ Krasznahorkay:2019lgi . This is also the most recent data for the proposed X17 boson resulting from the 8Be anomaly, regardless of the significant anomalies observed in the case of direct proton capture Sas:2022pgm .
In recent years, there were further studies that suggest potential connections between X17 boson and additional anomalies observed in the decays of excited 4He Firak:2020eil ; Krasznahorkay:2019lyl ; Krasznahorkay:2021joi and 12Krasznahorkay:2022pxs . These anomalies, like the 8Be case, exhibit deviations from the expected decay distributions. The masses of the proposed particle, as inferred from the anomalies observed in excited 4He and 12C decays, are reported as $16.94\pm 0.12({\rm stat})\pm 0.21({\rm syst})~{}\rm{MeV/c^{2}}$ Krasznahorkay:2021joi and $17.03\pm 0.11({\rm stat})\pm 0.2({\rm syst})~{}\rm{MeV/c^{2}}$ Krasznahorkay:2022pxs , respectively. It is remarkable that the mass values for the proposed particle as extracted from various experiments involving different nuclear systems are rather similar. This hints at a unique particle that can account for decay anomalies of excited 8Be, 4He, and 12C [^1] We note that there existed other experiments that perform searches for X17 boson as well. The search by VNU, Vietnam Anh:2024req confirmed the ${}^{8}{\rm Be}^{*}$ decay anomaly while the search by MEG II collaboration at PSI did not observe the resonance MEGII:2024urz . .
@@
where $q$ stands for $u$, $d$ and $s$ quarks, while $Q$ represents heavy quarks. From the data of ${}^{8}{\rm Be}$ anomaly, $R_{X}=5.8\times 10^{-6}$, with the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, it has been established that Feng:2016jff ; Fornal:2017msy $\varepsilon_{u}\simeq\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}\simeq\mp 7.4\times 10^{-3}$, where the null result of $\pi^{0}\to X\gamma$ search by NA48/2 NA482:2015wmo has been used. Further refined analysis takes into account isospin effects and their mixings Feng:2016ysn . Such an analysis considers three distinct isospin scenarios: no isospin effects, isospin mixing, and isospin mixing & breaking. Assuming that X17 is a vector boson, a statistical $\chi^{2}$ test based upon data from anomalous transitions observed in excited ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$ has been performed in Denton:2023gat . Utilizing eleven measurements that include angular and width data from the anomalous transitions in ATOMKI experiments, and the null result of $\pi^{0}\to X\gamma$ search NA482:2015wmo under the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, the above $\chi^{2}$ test determines the ranges of three parameters, $m_{X},\varepsilon_{n},\,{\rm and}\,\varepsilon_{p}$ where the latter two are related to the quark-level couplings by $\varepsilon_{p}=2\varepsilon_{u}+\varepsilon_{d}$ and $\varepsilon_{n}=2\varepsilon_{d}+\varepsilon_{u}$. It was found that  Denton:2023gat $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ and $\left|\varepsilon_{d}\right|\simeq(2.5-2.9)\times 10^{-3}$ with $\varepsilon_{u}$ and $\varepsilon_{d}$ in opposite signs. Specifically the most favored parameter set is $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ if isospin effects are not considered. With both isospin mixing and breaking effects considered, the most favored parameter set becomes $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$. It is noteworthy that these fitted values for $\left|\varepsilon_{u}\right|$ and $\left|\varepsilon_{d}\right|$ are much smaller than those determined earlier.
@@ II X17 contributions to heavy meson decays $H^{*}\to H+e^{+}+e^{-}$
For the case of $D^{*}$ meson decays, the form factors $F_{H^{*}H\gamma}(q)$ and $F_{H^{*}HX}(q)$ were calculated in Castro:2021gdf .
Figure 1: The experimentally measured as well as theoretically predicted values for $R_{D_{s}^{*+}}$. Various theoretical predictions are distinguished by different values taken for $\varepsilon_{c}$ and $\varepsilon_{s}$. VMD__\__ 2021 stands for the calculation of Castro:2021gdf using VMD approach and employing $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=7.4\times 10^{-3}$ extracted by Feng:2016jff ; Fornal:2017msy . VMD__\__ 2021 (corrected result) refers to our recalculation of $R_{D_{s}^{*+}}$ using the same values of $\varepsilon_{c}$ and $\varepsilon_{s}$. VMD__\__ 2023 denoted in blue represents our result of using $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for ${}^{8}{\rm Be}^{*}$ nuclear decays, and that denoted in black is obtained by employing $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered Denton:2023gat . The SM prediction is denoted as SM__\__ Photon.
We note that the values of $\varepsilon_{Q}$ and $\varepsilon_{q}$ significantly affect $R_{H^{*}}^{X}$. Fig. 1 presents the experimentally measured value of $R_{D^{*+}_{s}}$ as well as the corresponding theoretical predictions with and without contributions from the X17 boson. The measured value of $R_{D^{*+}_{s}}$ is marked in red while the SM prediction of this ratio is marked in yellow, which considers only the photon contribution and gives $R_{D^{*+}_{s}}^{\gamma}=6.7\times 10^{-3}$. This is consistent with the experimental data, $R_{D^{*+}_{s}}^{\text{data}}=\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ CLEO:2011mla . With the contribution of X17 boson included, VMD__\__ 2023 contains two cases. The first one, denoted in blue, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for ${}^{8}{\rm Be}^{*}$ nuclear decays. The second case, denoted in black, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered Denton:2023gat . The uncertainties for theoretical predictions arise from the uncertainty of coupling parameter $\lambda$ in the interaction vertex $\lambda H^{*}HV$ with $V=\rho^{0},\ \omega,\ \phi$. This parameter appears in the form factors $F_{H^{*}HX}(q^{2})$ and $F_{H^{*}H\gamma}(q^{2})$ and takes the value Colangelo:1993zq
@@
Theoretical predictions denoted by blue and black line segments in Fig. 1 are consistent with the experimental data. On the other hand, $R_{D_{s}^{*}}$ calculated with $\varepsilon_{u}=$$\pm 3.7\times 10^{-3},\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff yields $(3.3\pm 1.3)\times 10^{-2}$, which significantly deviates from the data. Our result disagrees with the calculation by Castro:2021gdf which gives $R_{D_{s}^{*+}}=(9.82\pm 0.60)\times 10^{-3}$. The latter turns out to be not far from the measurement $R_{D^{*+}_{s}}^{\text{data}}$.
Table 1: The values $R_{H^{*}}^{\gamma}$ and $R_{H^{*}}^{X}$ calculated with VMD model and the coupling values $\varepsilon_{u}=\varepsilon_{c}=\pm 5.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.9\times 10^{-3}$ that corresponds to best-fit parameters without isospin effects in ${}^{8}{\rm Be}^{*}$ decays considered Denton:2023gat .
@@
$D^{\ast 0}\rightarrow D^{0}e^{+}e^{-}$ $6.7\times 10^{-3}$ $5.5\times 10^{-7}$ $6.7\times 10^{-3}$ $(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq
$D^{\ast+}\rightarrow D^{+}e^{+}e^{-}$ $6.6\times 10^{-3}$ $(1.8\pm 0.1)\times 10^{-4}$ $6.8\times 10^{-3}$
$D_{s}^{\ast+}\rightarrow D_{s}^{+}e^{+}e^{-}$ $6.7\times 10^{-3}$ $(5.7\pm 3.0)\times 10^{-3}$ $(1.2\pm 0.3)\times 10^{-2}$ $(7.2_{-1.6}^{+1.8})\times 10^{-3}$ CLEO:2011mla
@@
We consider a similar analysis for the decay $D^{*0}\rightarrow D^{0}e^{+}e^{-}$. The experimental measurement gives $R^{\text{exp}}_{D^{*0}}=(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq and it is marked in red in Fig. 2. The SM_photon scenario yields $R_{D^{*0}}^{\gamma}=6.7\times 10^{-3}$. The full contributions $R_{D^{*0}}\equiv R_{D^{*0}}^{\gamma}+R_{D^{*0}}^{X}$ corresponding to different ranges of new physics couplings are labeled as VMD_2021 and VMD_2023, respectively. Table 1 presents results of $R_{H^{*}}$ arising from intermediate photon and X17 boson contributions for three $D^{*}$ meson decay modes within the VMD model, employing $\varepsilon_{u}=\varepsilon_{c}=\pm 5.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.9\times 10^{-3}$ that corresponds to best-fit parameters without isospin effects in ${}^{8}{\rm Be}^{*}$ decays considered Denton:2023gat . Considering isospin mixing and breaking effects in ${}^{8}{\rm Be}^{*}$ decays, i.e., $\varepsilon_{u}=\varepsilon_{c}=\pm 9.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.5\times 10^{-3}$, the corresponding $R_{H^{*}}$ values are $6.7\times 10^{-3}$, $6.8\times 10^{-3}$, and $(1.0\pm 0.4)\times 10^{-2}$ for $D^{*0}$, $D^{*+}$, and $D^{*+}_{s}$ decays, respectively.
@@
From Fig. 2, one can see that the observed $R_{D^{*0}}$ is much larger than every theoretical prediction. This implies that the intermediate X17 boson does not significantly enhance $D^{*0}\to D^{0}e^{+}e^{-}$ decay to match with the experimental data. The most straightforward remedy for the situation is by increasing the value of $\varepsilon_{c}$. However, an overly large $\varepsilon_{c}$ can easily disrupt the agreement between theory and experiment on $R_{D^{*}_{s}}$. To circumvent this problem, the value of $\varepsilon_{s}$ must also be adjusted accordingly. It is clear that we have to remove the assumption of $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$.

### v2

@@ Abstract
The recent ATOMKI experiments provided evidence pointing towards the existence of an X17 boson in the anomalous nuclear transitions of Beryllium-8, Helium-4, and Carbon-12. In this work, we consider X17 boson contributions to the previously measured $D$ meson decays which include $D_{s}^{*+}\rightarrow D_{s}^{+}e^{+}e^{-}$, $D_{s}^{*+}\rightarrow D_{s}^{+}\gamma$, $D^{*0}\rightarrow D^{0}e^{+}e^{-}$, and $D^{*0}\rightarrow D^{0}\gamma$, as well as the measured decays of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, and $\phi\rightarrow\eta\gamma$. Using the data of the above meson decays, we perform a fitting to the coupling parameters $\varepsilon_{u},\varepsilon_{c}$, and $\varepsilon_{s}$ by treating the couplings $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent from each other rather than assuming the generation universality $\varepsilon_{u}=\varepsilon_{c}$. It is found that the above fitting renders $|\varepsilon_{c}|=6.4\times 10^{-3}$, $|\varepsilon_{s}|=2.0\times 10^{-3}$ and a huge magnitude for $\varepsilon_{u}$, which is in serious tension with $\left|\varepsilon_{u}\right|$ determined from ATOMKI measurements.
@@ I Introduction
The results reported in Krasznahorkay:2015iga gives $m_{X}=16.70\pm 0.35({\rm stat})\pm 0.5({\rm syst})~{}\rm{MeV}$ and the ratio $\Gamma(^{8}{\rm Be}(18.15)\rightarrow~{}^{8}{\rm Be}+{X})\times{\rm BR}({X}\rightarrow e^{+}e^{-})/\Gamma(^{8}{\rm Be}(18.15)\rightarrow~{}^{8}{\rm Be}+\gamma)\equiv R_{X}=5.8\times 10^{-6}$, with a statistical significance of $6.8\sigma$. Later analysis from more 8Be experiments gave the updated results, $m_{X}=17.01\pm 0.16~{}\rm{MeV}$ and $R_{X}=(6\pm 1)\times 10^{-6}$ Krasznahorkay:2019lgi . This is also the most recent data for the proposed X17 boson resulting from the 8Be anomaly, regardless of the significant anomalies observed in the case of direct proton capture Sas:2022pgm .
In recent years, there were further studies that suggest potential connections between X17 boson and additional anomalies observed in the decays of excited 4He Firak:2020eil ; Krasznahorkay:2019lyl ; Krasznahorkay:2021joi and 12Krasznahorkay:2022pxs . These anomalies, like the 8Be case, exhibit deviations from the expected decay distributions. The masses of the proposed particle, as inferred from the anomalies observed in excited 4He and 12C decays, are reported as $16.94\pm 0.12({\rm stat})\pm 0.21({\rm syst})~{}\rm{MeV}$ Krasznahorkay:2021joi and $17.03\pm 0.11({\rm stat})\pm 0.2({\rm syst})~{}\rm{MeV}$ Krasznahorkay:2022pxs , respectively. It is remarkable that the mass values for the proposed particle as extracted from various experiments involving different nuclear systems are rather similar. This hints at a unique particle that can account for decay anomalies of excited 8Be, 4He, and 12C states [^1] We note that there existed other experiments that perform searches for X17 boson as well. The search by VNU, Vietnam Anh:2024req confirmed the ${}^{8}{\rm Be}^{*}$ decay anomaly while the search by MEG II collaboration at PSI did not observe the resonance MEGII:2024urz . .
@@
where $q$ stands for $u$, $d$ and $s$ quarks, while $Q$ represents heavy quarks. From the data of ${}^{8}{\rm Be}$ anomaly, $R_{X}=5.8\times 10^{-6}$, with the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, it has been established that Feng:2016jff ; Fornal:2017msy $\varepsilon_{u}\simeq\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}\simeq\mp 7.4\times 10^{-3}$, where the null result of $\pi^{0}\to X\gamma$ search by NA48/2 NA482:2015wmo has been used. Further refined analysis takes into account isospin effects and their mixings Feng:2016ysn . Such an analysis considers three distinct isospin scenarios: no isospin effects, isospin mixing, and isospin mixing & breaking. Assuming that X17 is a vector boson, a statistical $\chi^{2}$ test based upon data from anomalous transitions observed in excited ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$ has been performed in Denton:2023gat . Utilizing eleven measurements that include angular and width data from the anomalous transitions in ATOMKI experiments, and the null result of $\pi^{0}\to X\gamma$ search NA482:2015wmo under the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, the above $\chi^{2}$ test determines the ranges of three parameters, $m_{X},\varepsilon_{n},\,{\rm and}\,\varepsilon_{p}$ where the latter two are related to the quark-level couplings by $\varepsilon_{p}=2\varepsilon_{u}+\varepsilon_{d}$ and $\varepsilon_{n}=2\varepsilon_{d}+\varepsilon_{u}$. It was found that  Denton:2023gat $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ and $\left|\varepsilon_{d}\right|\simeq(2.5-2.9)\times 10^{-3}$ with $\varepsilon_{u}$ and $\varepsilon_{d}$ in opposite signs. Specifically the most favored parameter set is $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ and $m_{X}=16.84\ {\rm MeV}$ if isospin effects are not considered. With both isospin mixing and breaking effects considered, the most favored parameter set becomes $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ and $m_{X}=16.85\ {\rm MeV}$. It is noteworthy that the above fitted values for $\left|\varepsilon_{u}\right|$ and $\left|\varepsilon_{d}\right|$ are much smaller than those determined earlier.
@@ II X17 contributions to heavy meson decays $H^{*}\to H+e^{+}+e^{-}$
For the case of $D^{*}$ meson decays, the form factors $F_{H^{*}H\gamma}(q^{2})$ and $F_{H^{*}HX}(q^{2})$ were parameterized as follows Castro:2021gdf :
$F_{H^{*}H\gamma}(q^{2})$ $=$ $\sqrt{\frac{m_{H^{*}}}{m_{H}}}\left(\frac{e_{Q}}{m_{H^{*}}}+\frac{e_{q}}{m_{q}(q^{2})}\right),$
$F_{H^{*}HX}(q^{2})$ $=$ $\sqrt{\frac{m_{H^{*}}}{m_{H}}}\left(\frac{\varepsilon_{Q}}{m_{H^{*}}}+\frac{\varepsilon_{q}}{m_{q}(q^{2})}\right),$ (9)
where the effective light-quark mass parameter $m_{q}(q^{2})$ can be calculated by VMD Colangelo:1993zq .
Figure 1: The experimentally measured as well as theoretically predicted values for $R_{D_{s}^{*+}}$. Various theoretical predictions are distinguished by different values taken for $\varepsilon_{c}$ and $\varepsilon_{s}$. VMDa stands for the calculation of Castro:2021gdf using VMD approach and employing $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$ extracted by Feng:2016jff ; Fornal:2017msy . VMDa Corrected Result refers to our recalculation of $R_{D_{s}^{*+}}$ using the same values of $\varepsilon_{c}$ and $\varepsilon_{s}$. VMDb denoted in blue represents our result of using $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for ${}^{8}{\rm Be}^{*}$ nuclear decays, and that denoted in black is obtained by employing $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered Denton:2023gat . The SM prediction is denoted as SM__\__ Photon.
Inferring from Eqs. (27) and (28) of Ref. Colangelo:1993zq , we obtain
$m_{u,d}(q^{2})^{-1}$ $=$ $-\sum_{V=\rho^{0},\omega}\left(\sqrt{2}g_{V}\lambda\frac{f_{V}}{m_{V}^{2}}\right)\left(1-\frac{q^{2}}{m_{V}^{2}}\right)^{-1}$
$m_{s}(q^{2})^{-1}$ $=$ $-\left(2\sqrt{2}g_{V}\lambda\frac{f_{\phi}}{m_{\phi}^{2}}\right)\left(1-\frac{q^{2}}{m_{\phi}^{2}}\right)^{-1},$ (10)
with $\lambda$ the strength of heavy-meson coupling to light vector mesons described by the Lagrangian
[…]
$\langle 0|\frac{1}{\sqrt{2}}(\bar{u}\gamma_{\mu}u-\bar{d}\gamma_{\mu}d)|\rho^{0}(\epsilon)\rangle$ $=$ $\epsilon_{\mu}f_{\rho^{0}},$
$\langle 0|\frac{1}{\sqrt{2}}(\bar{u}\gamma_{\mu}u+\bar{d}\gamma_{\mu}d)|\omega(\epsilon)\rangle$ $=$ $\epsilon_{\mu}f_{\omega},$
$\langle 0|\bar{s}\gamma_{\mu}s|\phi(\epsilon)\rangle$ $=$ $\epsilon_{\mu}f_{\phi},$ (12)
where $\epsilon_{\mu}$ is the polarization of the vector meson. Using the current data of $V\to e^{+}e^{-}$ ParticleDataGroup:2024cfk , it is found that $(f_{\rho^{0}},f_{\omega},f_{\phi})=(0.171,0.155,0.232)\ {\rm GeV}^{2}$ Castro:2021gdf . Finally the parameter $\lambda$ can be determined by the data of semi-leptonic decays $D\to\pi(K)l\nu_{l}$ and $D\to K^{*}l\nu_{l}$ as pointed out in Colangelo:1993zq . With the updated data, it is found that Castro:2021gdf
@@
We point out that our $m_{s}(q^{2})^{-1}$ agrees with that of Castro:2021gdf . On the other hand, $m_{u,d}(q^{2})^{-1}$ given by Eq. (5) of Castro:2021gdf is twice larger than ours. It is easily seen that our results in Eq. (10) satisfy SU(3) symmetry relation $m_{u}(q^{2})^{-1}=m_{d}(q^{2})^{-1}=m_{s}(q^{2})^{-1}$ in the SU(3) symmetry limit $f_{\rho^{0}}=f_{\omega}=f_{\phi}$, and $m_{\rho^{0}}=m_{\omega}=m_{\phi}$.
Fig. 1 presents the experimentally measured value of $R_{D^{*+}_{s}}$ as well as the corresponding theoretical predictions with and without contributions from the X17 boson. The measured value of $R_{D^{*+}_{s}}$ is marked in red while the SM prediction of this ratio is marked in yellow, which considers only the photon contribution and gives $R_{D^{*+}_{s}}^{\gamma}=6.7\times 10^{-3}$. This is consistent with the experimental data, $R_{D^{*+}_{s}}^{\text{exp}}=\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ CLEO:2011mla . Including the contribution by X17 boson with $\varepsilon_{u}=$$\pm 3.7\times 10^{-3},\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff , we obtain $R_{D^{*+}_{s}}=(2.6\pm 1.4)\times 10^{-2}$ as denoted by “VMDa Corrected Result”, which is in a mild tension with the data. For current and subsequent calculations, we take $m_{X}=17$ MeV. It is found that the decay ratios, such as $R_{D^{*+}_{s}}$ and those involving other mesons, are not sensitive to $m_{X}$ over its allowed range. Our result disagrees with the calculation by Castro:2021gdf which gives $R_{D_{s}^{*+}}=(9.82\pm 0.60)\times 10^{-3}$. The detailed comparison between two results will be given in the next paragraph. The results denoted by VMDb are obtained with $\varepsilon_{u,d}$ taken from Denton:2023gat . They are divided into two cases. The first case, denoted in blue, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for ${}^{8}{\rm Be}^{*}$ nuclear decays. The second case, denoted in black, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered. The uncertainties for theoretical predictions arise from the uncertainty of coupling parameter $\lambda$. It is seen that both predictions under VMDb are consistent with the experimental data.
To compare with Ref. Castro:2021gdf in details, we note that our calculations reproduce $\varepsilon_{Q}/m_{H^{*}}=1.75\times 10^{-3}\ {\rm GeV}^{-1}$ and $\varepsilon_{q}/m_{q}(m_{X}^{2})=-7.83\times 10^{-3}\ {\rm GeV}^{-1}$ for the decay $D_{s}^{*+}\to D_{s}^{+}X$ given in Table II of that paper. Hence $F_{H^{*}HX}(m_{X}^{2})$ defined by Eq. (9), is easily obtained as $(-0.63\pm 0.05)\times 10^{-2}\ {\rm GeV}^{-1}$ rather than $(-0.91\pm 0.06)\times 10^{-2}\ {\rm GeV}^{-1}$ given in Table II of Castro:2021gdf . Here the uncertainty of $F_{H^{*}HX}(m_{X}^{2})$ is due to the uncertainty of $\lambda$. Furthermore, the quantity $e_{q}/m_{q}(0)$ in $D_{s}^{*+}\to D_{s}^{+}\gamma$ is related to $\varepsilon_{q}/m_{q}(m_{X}^{2})$ by a simple scaling: $e_{q}/m_{q}(0)=e_{q}/\varepsilon_{q}\cdot\varepsilon_{q}/m_{q}(0)\simeq e_{q}/\varepsilon_{q}\cdot\varepsilon_{q}/m_{q}(m_{X}^{2})$. With $e_{s}=-1/3$, $\varepsilon_{s}=-7.4\times 10^{-3}$, and $\varepsilon_{s}/m_{s}(m_{X}^{2})=-7.83\times 10^{-3}\ {\rm GeV}^{-1}$ as just mentioned, one obtains $e_{s}/m_{s}(0)=-0.35\ {\rm GeV}^{-1}$ rather than $-0.48\ {\rm GeV}^{-1}$ given in Table I of  Castro:2021gdf . Since $e_{c}/m_{D^{*+}_{s}}=0.32\ {\rm GeV}^{-1}$, one obtains $F_{D^{*+}_{s}D^{+}_{s}\gamma}(0)=-0.038\pm 0.020\ {\rm GeV}^{-1}$ where the uncertainty again arises from the uncertainty of $\lambda$. We observe that the central value of $F_{D^{*+}_{s}D^{+}_{s}\gamma}(0)$ is much smaller than that in Castro:2021gdf which gives $F_{D^{*+}_{s}D^{+}_{s}\gamma}(0)=-0.17\pm 0.03\ {\rm GeV}^{-1}$. Hence we predict a much larger $R^{X}_{D^{*+}_{s}}$ due to a much smaller $F^{2}_{D^{*+}_{s}D^{+}_{s}\gamma}(0)$ on the right hand side of Eq. (7). This leads to a much larger overall ratio $R_{D^{*+}_{s}}$.
Table 1: The values $R_{H^{*}}^{\gamma}$ and $R_{H^{*}}^{X}$ calculated with VMD model and the coupling values $\varepsilon_{u}=\varepsilon_{c}=\pm 5.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.9\times 10^{-3}$ that corresponds to best-fit parameters without considering isospin effects in ${}^{8}{\rm Be}^{*}$ decays Denton:2023gat .
@@
$D^{\ast 0}\rightarrow D^{0}e^{+}e^{-}$ $6.7\times 10^{-3}$ $5.6\times 10^{-7}$ $6.7\times 10^{-3}$ $(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq
$D^{\ast+}\rightarrow D^{+}e^{+}e^{-}$ $6.6\times 10^{-3}$ $(1.2\pm 0.4)\times 10^{-3}$ $(7.8\pm 0.4)\times 10^{-3}$
$D_{s}^{\ast+}\rightarrow D_{s}^{+}e^{+}e^{-}$ $6.7\times 10^{-3}$ $(4.2\pm 3.1)\times 10^{-3}$ $(1.1\pm 0.3)\times 10^{-2}$ $(7.2_{-1.6}^{+1.8})\times 10^{-3}$ CLEO:2011mla
@@
We consider a similar analysis for the decay $D^{*0}\rightarrow D^{0}e^{+}e^{-}$. The experimental measurement gives $R^{\text{exp}}_{D^{*0}}=(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq and it is marked in red in Fig. 2. The SM contribution due to the photon mediation yields $R_{D^{*0}}^{\gamma}=6.7\times 10^{-3}$. The full contributions $R_{D^{*0}}\equiv R_{D^{*0}}^{\gamma}+R_{D^{*0}}^{X}$ corresponding to different ranges of new physics couplings are labeled as VMDa and VMDb, respectively. Table 1 presents results of $R_{H^{*}}$ arising from intermediate photon and X17 boson contributions for three $D^{*}$ meson decay modes within the VMD model, employing $\varepsilon_{u}=\varepsilon_{c}=\pm 5.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.9\times 10^{-3}$ that corresponds to best-fit parameters without isospin effects in ${}^{8}{\rm Be}^{*}$ decays considered Denton:2023gat . Considering isospin mixing and breaking effects in ${}^{8}{\rm Be}^{*}$ decays, i.e., $\varepsilon_{u}=\varepsilon_{c}=\pm 9.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.5\times 10^{-3}$, the corresponding $R_{H^{*}}$ values are $6.7\times 10^{-3}$, $(7.4\pm 0.3)\times 10^{-3}$, and $(9.3\pm 1.9)\times 10^{-3}$ for $D^{*0}$, $D^{*+}$, and $D^{*+}_{s}$ decays, respectively.
@@
From Fig. 2, one can see that the observed $R_{D^{*0}}$ is much larger than every theoretical prediction. This implies that the intermediate X17 boson does not significantly enhance $D^{*0}\to D^{0}e^{+}e^{-}$ decay to match with the experimental data. The most straightforward remedy for the problem is by increasing the value of $\varepsilon_{c}$. However, an overly large $\varepsilon_{c}$ can easily disrupt the agreement between theory and experiment on $R_{D^{*}_{s}}$. To circumvent this problem, the value of $\varepsilon_{s}$ must also be adjusted accordingly. It is clear that we have to remove the assumption of $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$.